OpenStudy (anonymous):

Can anyone help me with the following question? 3-3^x=5^x. Find the value x. Thank you.

OpenStudy (turingtest):

\[3-3^x=5^x\]and you want to solve for x ?

OpenStudy (anonymous):

Yes.

OpenStudy (turingtest):

I'm sorry, I know of no analytic way to do this http://www.wolframalpha.com/input/?i=solve%203-3%5Ex%3D5%5Ex&t=crmtb01

OpenStudy (asnaseer):

have you been asked to use any particular method? e.g. iteration, Newton-Raphson, etc...

OpenStudy (anonymous):

no. I thought can use the logarithm rules to solve it. Is it possible?

OpenStudy (asnaseer):

no - not with logs alone as you would end up with:\[\log(3-3^x)=\log(5^x)=x\log(5)\]and there is no way to simplify \(\log(3-3^x)\). The only way I know uses iterations to converge to the value.

OpenStudy (asnaseer):

A simple iterative way to solve this would be to rearrange it as follows:\[x_{n+1}=\frac{\log(3-3^{x_n})}{\log(5)}\]then start with, say, \(x_0=0\) and work out \(x_1\). plug this back in to work out \(x_2\) and keep going until you converge to a solution.

OpenStudy (anonymous):

ok then. Thank you very much.

OpenStudy (asnaseer):

yw

OpenStudy (anonymous):

A numerical approximation is x=0.297324